OpenStudy (anonymous):

Assume the y varies inversely with x and y = 5 when x = 1.3. Write the indirect variation equation for the relationship. would I just multiply 5*1.3?

OpenStudy (anonymous):

y=k/x

OpenStudy (anonymous):

how do I know what k is or what am I doin?

OpenStudy (anonymous):

k is just a constant, it could be any number but its just there in case (to be more mathematical). y=1/x shows the relationship too, for what information is given. you can just plug in y and x to y=1/x

OpenStudy (anonymous):

because a fraction with a larger denominator is smaller, as x grows y gets smaller in y=1/x. This is what they mean by indirect variation.

jimthompson5910 (jim_thompson5910):

y=k/x y=k/1.3 ... plug in x = 1.3 5 = k/1.3 ... plug in y = 5 now solve for k

OpenStudy (anonymous):

o okay just a sec. and ill tel you what I get

OpenStudy (anonymous):

one of my chooses is 3.8 that's the only one that could work out of the chooses I got

jimthompson5910 (jim_thompson5910):

5 = k/1.3 5*1.3 = k ... Multiply both sides by 1.3 to isolate k. 6.5 = k k = 6.5

jimthompson5910 (jim_thompson5910):

so the equation is y = 6.5/x

OpenStudy (anonymous):

o sorry I did something totally different but I do understand what to do now thank you very much!

jimthompson5910 (jim_thompson5910):

ok great, glad you know what to do now